OpenStudy (anonymous):

4/3 of a number is 22 what is 8/3 of that number

OpenStudy (raden):

just multiply by 2

OpenStudy (shamim):

let the number is x

OpenStudy (shamim):

so 4x/3=22

OpenStudy (shamim):

x=22*3/4

OpenStudy (shamim):

x=33/2

OpenStudy (shamim):

so ur new number will b (8/3)*(33/2)

OpenStudy (shamim):

44

OpenStudy (raden):

(4/3)x = 22 times 2 both sides (8/3)x = 44 the result is same ^